# Dents in the Mirror: A Novel Probe of Dark Matter Substructure in Galaxy Clusters from the Astrometric Asymmetry of Lensed Arcs

### v1

@@ Abstract
Astrometric perturbations of lensed arcs behind galaxy clusters have been recently suggested as promising probes of small-scale ($\lesssim 10^{9}M_{\odot}$) dark matter substructure. Populations of cold dark matter (CDM) subhalos, predicted in hierarchical structure formation theory, can break the symmetry of arcs near the critical curve, leading to positional shifts in the observed images. We present a novel statistical method to constrain the average subhalo mass fraction ($f_{\rm sub}$) in clusters that takes advantage of this induced positional asymmetry. Focusing on CDM, we extend a recent semi-analytic model of subhalo tidal evolution to accurately simulate realistic subhalos within a cluster-scale host. We simulate the asymmetry of lensed arcs from these subhalo populations using Approximate Bayesian Computation. Using mock data, we demonstrate that our method can reliably recover the simulated $f_{\rm sub}$ to within 68% CI in 73% of cases, regardless of the lens model, astrometric precision, and image morphology. We show that the constraining power of our method is optimized for larger samples of well observed arcs, ideal for recent JWST observations of cluster lenses. As a preliminary test, we apply our method to the MACSJ0416 Warhol arc and AS1063 System 1. For Warhol we constrain the upper limit on $\log f_{\rm sub}<-3.40^{+1.06}_{-0.97}$, while for AS1063 System 1 we constrain $\log f_{\rm sub}=-2.36^{+0.56}_{-0.89}$ (both at 68% CI), consistent with CDM predictions. We elaborate on our method’s limitations and its future potential to place stringent constraints on dark matter properties in cluster environments.
@@
pubyear: 2025pagerange: Dents in the Mirror: A Novel Probe of Dark Matter Substructure in Galaxy Clusters from the Astrometric Asymmetry of Lensed ArcsA
@@ 1 Introduction
Current observations of the cosmic microwave background strongly support the theory that dark matter is cold and collisionless (planck20). In this standard cosmological model, cold dark matter (CDM) halos form hierarchically at all length scales of cosmological significance (davis85; klypin99; moore99) and their density profiles can be universally described by the Navarro-Frenk-White (NFW) profile (nfw97). Additionally, a universal subhalo mass function (SHMF) is expected to describe CDM halos at all scales (giocoli08). This is particularly useful in studying small-scale CDM substructures within a main host halo, hereafter referred to as subhalos. CDM subhalos are predicted in the standard CDM paradigm, however this paradigm fails to explain some small-scale observations of galaxies (see for review e.g. delpopolo17; bullock17; sales22). As a result, a new frontier for astrophysical probes of dark matter is the search for populations of small-scale ($\lesssim 10^{9}M_{\odot}$) CDM subhalos, the detection of which can help to determine the true nature of dark matter.
@@
One such unique probe of dark matter is strong gravitational lensing, which is ideal due to its sensitivity to the gravitational potential of mass structures at all scales. Using galaxy-scale lenses, dark matter substructures have been prolifically studied, with numerous constraints having been made on dark matter properties using subhalos (dalal02; vegetti14; despali17; birrer17; hsueh20; gilman20; he22; gilman24; keeley24). For galaxy cluster-scale lenses, dark matter constraints from subhalos are much more sparse, mostly restricted to intermediate mass ($\sim 10^{11}M_{\odot})$ halos (natarajan04; natarajan17) and the distribution of subhalos (umetsu16). Recently, the discovery of numerous microlensed individual stars with the Hubble Space Telescope (HST) (kelly22) and James Webb Space Telescope (JWST) (windhorst23; yan23; fudamoto2025) have extended the search for dark matter subhalos in galaxy clusters by providing high resolution observational data on sub-arcsecond angular scales. As a result, new methods to probe subhalos or other small-scale dark matter structures in galaxy clusters have been recently developed (venumadhav17; dai18; dai20a; dai20c; williams24; broadhurst25; palencia25). In particular, optically faint or invisible CDM subhalos may be responsible for the puzzling aspects of some high-$z$ magnified sources (diego22b; ji25; pascale25).
@@
Many of these new methods focus on CDM subhalo perturbations to the cluster’s critical curve. Given that the critical curve indicates the region of highest magnification in the gravitational lens, the observational signatures of these CDM subhalos are primarily flux ratio anomalies within the images (dai20a; ji25). This is a similar concept as is done on galaxy scales with gravitational imaging (koopmans05; vegetti09), a method that has been successful at constraining the CDM subhalo SHMF (vegetti18; ritondale19). At cluster-scales, an underappreciated effect of these perturbations is the astrometric shifts of the lensed image positions (dai18; abe23). Any detection of these positional shifts of images would be evidence ruling out a smooth density profile on small scales, as is commonly implied from lens models (limousin22). The difficulty lies in disentangling between astrometric uncertainty from the observed image positions and a genuine perturbation. Additionally, disentangling whether a more complex lens model can account for an observed image shift, rather than a local perturbation, poses another challenge. The former can be addressed with new high resolution data from JWST, while the latter can be accounted for with a statistical or model agnostic approach.
@@
An important and necessary prerequisite for these types of studies is the development of sophisticated models of CDM subhalos. The physical interaction of CDM subhalos within their host leads to strong tidal stripping effects (hayashi03). This tidal evolution is dependent on the host halo mass and the trajectory of the orbiting subhalo, thus making models of the process statistical in nature (kravtsov04; han16). This has motivated the creation of high resolution cosmological simulations of CDM to create precise theoretical predictions of subhalo density structure (Springel2008). These predictions allow for an ease of use of physically realistic CDM subhalo structure that can be quickly applied to a wide variety of dark matter probes (taylor01; penarrubia08; benson12; du24).
@@
In this paper, we present a novel method to investigate the astrometric shifts of images in cluster lenses. The fundamental goal of this work is to constrain the CDM subhalo SHMF based on the induced perturbations from populations of CDM subhalos near the lensing critical curve. To ease comparison with previous works, we focus on the mean subhalo mass fraction ($f_{\rm sub}$) as the primary parameter of interest, implicitly assuming that it is the same in both galaxy and cluster environments. This parameter can be directly calculated from the subhalo SHMF, and has been constrained in many studies using galaxy-scale lenses (vegetti14; despali17; hsueh20; gilman20), but not with cluster-scale lenses. For our method, we utilize a known result from gravitational lensing in that sources that form near a fold caustic will form symmetric image pairs across the critical curve (schneider92). Thus, astrometric perturbations from a population of CDM subhalos lying in the lens plane will manifest as an asymmetry [^1] The induced asymmetry from CDM subhalos is akin to “Denting the mirror”, hence the title. in the image positions across the critical curve (dai18). We develop a methodology to infer the hidden subhalo population using a likelihood-free inference method that uses the asymmetry of lensed image pairs as a summary statistic. After validating the modeling framework on simulated datasets, we apply it to two lensed arcs in MACSJ0416 and AS1063 to infer the projected mass fraction in dark matter subhalos. In the future, the method demonstrated in this paper will be applied to larger samples of gravitational lenses to provide stringent constraints on $f_{\rm sub}$ from cluster lenses.
@@ 2.1 Gravitational Lensing Formalism
Here, we briefly review the strong gravitational lensing formalism necessary for this article. We refer the reader to various review articles (e.g blandford86; schneider92; narayan96) for additional details.
@@ 2.2 Lensing Near the Critical Curve
Cluster lens modeling analyses often model extended lensed arcs using identified counterimaged “knots” within the arc as constraints. Modeling the relative positions of lensed knots improves the cluster lens model by constraining the deflection field on angular scales comparable to the image separation (e.g. bergamini23; perera24b). These knots likely correspond to bright sub-galactic structures within the source galaxy (e.g. HII regions in the case of MACSJ1149 hwilliams24). Due to these knots forming on sub-arcsecond scales, they are treated as point images with lower positional uncertainty than a typical cluster scale image. For this paper, we represent lensed arcs as a collection of lensed point image knots across a critical curve.
@@
The prediction that image pairs collect along a straight line is a well known result in gravitational lensing near a fold caustic (blandford86; schneider92). We derive this property below. We begin with the lensing Jacobian as defined in equation 5. It is more useful to write the Jacobian in terms of the convergence and shear tensor components ($\gamma_{1}$ and $\gamma_{2}$):
@@
Usefully, $\mathcal{A}(\boldsymbol{\theta})=\partial\boldsymbol{\beta}/\partial\boldsymbol{\theta}$. Since we are close to the critical curve, we can use the same approximations as before, namely that the origin is located at the critical curve position ($\boldsymbol{\theta_{c}}=0$). We define the components of the displacement $\delta\boldsymbol{\theta}=(\delta\theta_{1},\delta\theta_{2})$ and source position $\boldsymbol{\beta}=(\beta_{1},\beta_{2})$. The subscripts 1 and 2 indicate the tangential and critical directions in the principal coordinate system. The leading order terms can be written as:
@@
Since $\delta\theta_{2}$ is the critical direction (perpendicular to the critical curve), this shows that the formed images near the caustic fold are in fact symmetric on either side. As a consequence, the midpoints of these images will form at the critical curve.
@@ 3.1.1 Cluster Lens Profile
To demonstrate the methodology presented in this work, we create a catalog of simulated cluster lens systems. These systems are comprised of a macro cluster lens profile, plus subhalos. We model the macro cluster-scale lens profile utilizing 3 main halos at a lens redshift $z_{d}=0.25$. Two of the halos are separated by 0.5" (1.97 kpc) such that their mass profile is effectively combined. This is done to mimic a slightly non-elliptical profile. The third profile is located a distance of 17.5" (68.8 kpc) from the other combined halos. Each halo is modelled as a Non-Singular Isothermal Ellipsoid (NSIE), which has an analytical lens potential (hinshaw87; halkola06):
@@
where $\sigma_{v}$ is the velocity dispersion, $R_{c}$ is the angular core radius [^2] The physical core radius for the NSIE is equal to $D_{d}R_{c}$., $q$ is the axis ratio, and the vector position $\boldsymbol{\theta}$ has components $(\theta_{1},\theta_{2})$. The total mass of the cluster-scale lens profile is $1.3\times 10^{14}M_{\odot}$. Large clusters such as this are the most efficient lenses akin to the Frontier Fields (lotz17), motivating our lens profile as reasonable and realistic. Properties of each NSIE halo are presented in Table 1. Figure 2 shows the surface mass density of the cluster lens. From here, we refer to this profile as the fiducial macrolens.
@@ 3.1.2 Lensed Arcs Near Caustic Folds
In Section 2.2 we described the general properties of lensed images near the critical curve. We also showed that sources that form near the caustic folds of cluster-scale gravitational lenses will appear as fully symmetric lensed arcs across a linear critical curve for a smooth lens profile (blandford86; schneider92; dai18). Here, we focus specifically on the corresponding approximations that we adopt.
@@
The addition of dark matter substructures in the region within $\sim 1-2$" of the critical curve will displace the smooth critical curve and can produce asymmetries in the arcs. It should be noted that the subhalo mass fraction is constrained to be on the order of $\sim 1\%$ (gilman19; dai20a; ji25), but these constraints come mainly from galaxy-scale lens systems. The addition of dark matter subhalos will not cause lensing perturbations on scales beyond the $\sim 1-2$" window. Therefore, we only simulate subhalos within the immediate vicinity of lensed arcs.
@@
According to gravitational lensing theory, the midpoints of images that form near a smooth critical curve lie along the critical curve (blandford86; venumadhav17; dai18) as we showed in Section 2.2. This symmetric setup can be broken with the inclusion of dark matter subhalos near the critical curve, as shown in Figure 1. Recall that images in these cases where the source is very near to the caustic manifest as extended arcs and are instead represented as lensed point image knots. Because of this, the image midpoints are easily calculated as the midpoints of the counterimaged knots. We note that this scenario where the midpoints lie along the critical curve is restricted to sources that form near to the caustic, and is not generally true for all source positions.
@@
When considering angular scales significantly smaller than the radius of curvature of the macrolens near the caustic fold, the forward lens-projected (i.e. mapped from the source to lens plane) critical curve can be approximated as a straight line, and we expect unperturbed image midpoints to collect along this straight line. In real examples of arcs, this appears to be relatively common (e.g. kaurov19; chen19; diego23c). In fact, this is an intuitively expected setup for lensed arcs, given that in clusters the caustic fold spans a significantly larger region than cusps. It should also be noted that this assumption regarding the image midpoints is commonly used as a proxy for the true location of the smooth critical curve (kelly22; broadhurst25), emphasizing the utility of this approximation.
@@
It is important to note that there can be a degeneracy when evaluating lensed arcs along the critical curve. Deviations from a smooth fold can occur due to the presence of galaxy scale perturbers, such as cluster member galaxies (e.g. dai20a). In this case, the midpoints will still lie along the critical curve, but the critical curve will no longer be a straight line. Likewise, large subhalos ($\gtrsim 10^{10}M_{\odot}$) may also contribute, albeit rarely, to the variance in the critical curve linearity. Furthermore, sources forming slightly away from the caustic fold will produce images across an intrinsically curved smooth critical curve, thus not allowing the straight line critical curve approximation. As such, there exists a degeneracy between deviations from the symmetry of arcs stemming from either dark matter subhalos or lens model complexities. For this work, we restrict the analysis to lensed arcs that are unaffected by cluster members or large subhalos and are subject to only the smooth macrolens near a caustic fold. It turns out that this need not be a strict requirement, but we caution anyway that our results should only be applied to lensed arcs that are consistent with this description. We examine in greater detail the limits of this assumption in Appendix A. In future works, these contributions away from linearity should be included in the statistical analysis.

### v2

@@ Abstract
Astrometric perturbations of lensed arcs behind galaxy clusters have been recently suggested as promising probes of small-scale ($\lesssim 10^{9}M_{\odot}$) dark matter substructure. Populations of cold dark matter (CDM) subhalos, predicted in hierarchical structure formation theory, can break the symmetry of arcs near the critical curve, leading to positional shifts in the observed images. We present a novel statistical method to constrain the average subhalo mass fraction ($f_{\rm sub}$) in clusters that takes advantage of this induced positional asymmetry. Focusing on CDM, we extend a recent semi-analytic model of subhalo tidal evolution to accurately simulate realistic subhalos within a cluster-scale host. We simulate the asymmetry of lensed arcs from these subhalo populations using Approximate Bayesian Computation. Using mock data, we demonstrate that our method can reliably recover the simulated $f_{\rm sub}$ to within 68% CI in 73% of cases, regardless of the lens model, astrometric precision, and image morphology. We show that the constraining power of our method is optimized for larger samples of well observed arcs, ideal for recent JWST observations of cluster lenses. As a preliminary test, we apply our method to the MACSJ0416 Warhol arc and AS1063 System 1. For Warhol we constrain the upper limit on $\log f_{\rm sub}<-3.48^{+1.00}_{-0.91}$, while for AS1063 System 1 we constrain $\log f_{\rm sub}=-2.44^{+0.61}_{-0.86}$ (both at 68% CI), consistent with CDM predictions. We elaborate on our method’s limitations and its future potential to place stringent constraints on dark matter properties in cluster environments.
@@
pubyear: 2025pagerange: Dents in the Mirror: A Novel Probe of Dark Matter Substructure in Galaxy Clusters from the Astrometric Asymmetry of Lensed ArcsD
@@ 1 Introduction
Current observations of the cosmic microwave background strongly support the theory that dark matter is cold and collisionless (Planck Collaboration et al., 2020). In this standard cosmological model, cold dark matter (CDM) halos form hierarchically at all length scales of cosmological significance (Davis et al., 1985; Klypin et al., 1999; Moore et al., 1999) and their density profiles can be universally described by the Navarro-Frenk-White (NFW) profile (Navarro et al., 1997). Additionally, a universal subhalo mass function (SHMF) is expected to describe CDM halos at all scales (Giocoli et al., 2008). This is particularly useful in studying small-scale CDM substructures within a main host halo, hereafter referred to as subhalos. CDM subhalos are predicted in the standard CDM paradigm, however this paradigm fails to explain some small-scale observations of galaxies (see for review e.g. Del Popolo and Le Delliou, 2017; Bullock and Boylan-Kolchin, 2017; Sales et al., 2022). As a result, a new frontier for astrophysical probes of dark matter is the search for populations of small-scale ($\lesssim 10^{9}M_{\odot}$) CDM subhalos, the detection of which can help to determine the true nature of dark matter.
@@
One such unique probe of dark matter is strong gravitational lensing, which is ideal due to its sensitivity to the gravitational potential of mass structures at all scales. Using galaxy-scale lenses, dark matter substructures have been prolifically studied, with numerous constraints having been made on dark matter properties using subhalos (Dalal and Kochanek, 2002; Vegetti et al., 2014; Despali and Vegetti, 2017; Birrer et al., 2017; Hsueh et al., 2020; Gilman et al., 2020; He et al., 2022; Gilman et al., 2024; Keeley et al., 2024). For galaxy cluster-scale lenses, dark matter constraints from subhalos are much more sparse, mostly restricted to intermediate mass ($\sim 10^{11}M_{\odot})$ halos (Natarajan and Springel, 2004; Natarajan et al., 2017) and the distribution of subhalos (Umetsu et al., 2016). Recently, the discovery of numerous microlensed individual stars with the Hubble Space Telescope (HST) (Kelly et al., 2022) and James Webb Space Telescope (JWST) (Windhorst et al., 2023; Yan et al., 2023; Fudamoto et al., 2025) have extended the search for dark matter subhalos in galaxy clusters by providing high resolution observational data on sub-arcsecond angular scales. As a result, new methods to probe subhalos or other small-scale dark matter structures in galaxy clusters have been recently developed (Venumadhav et al., 2017; Dai et al., 2018, 2020; Dai and Miralda-Escudé, 2020; Williams et al., 2024b; Broadhurst et al., 2025; Palencia et al., 2025). In particular, optically faint or invisible CDM subhalos may be responsible for the puzzling aspects of some high-$z$ magnified sources (Diego et al., 2022; Ji and Dai, 2025; Pascale et al., 2025).
@@
Many of these new methods focus on CDM subhalo perturbations to the cluster’s critical curve. Given that the critical curve indicates the region of highest magnification in the gravitational lens, the observational signatures of these CDM subhalos are primarily flux ratio anomalies within the images (Dai et al., 2020; Ji and Dai, 2025). This is a similar concept as is done on galaxy scales with gravitational imaging (Koopmans, 2005; Vegetti and Koopmans, 2009), a method that has been successful at constraining the CDM subhalo SHMF (Vegetti et al., 2018; Ritondale et al., 2019). At cluster-scales, an underappreciated effect of these perturbations is the astrometric shifts of the lensed image positions (Dai et al., 2018; Abe et al., 2024). Any detection of these positional shifts of images would be evidence ruling out a smooth density profile on small scales, as is commonly implied from lens models (Limousin et al., 2022). The difficulty lies in disentangling between astrometric uncertainty from the observed image positions and a genuine perturbation. Additionally, disentangling whether a more complex lens model can account for an observed image shift, rather than a local perturbation, poses another challenge. The former can be addressed with new high resolution data from JWST, while the latter can be accounted for with a statistical or model agnostic approach.
@@
An important and necessary prerequisite for these types of studies is the development of sophisticated models of CDM subhalos. The physical interaction of CDM subhalos within their host leads to strong tidal stripping effects (Hayashi et al., 2003). This tidal evolution is dependent on the host halo mass and the trajectory of the orbiting subhalo, thus making models of the process statistical in nature (Kravtsov et al., 2004; Han et al., 2016). This has motivated the creation of high resolution cosmological simulations of CDM to create precise theoretical predictions of subhalo density structure (Springel et al., 2008). These predictions allow for an ease of use of physically realistic CDM subhalo structure that can be quickly applied to a wide variety of dark matter probes (Taylor and Babul, 2001; Peñarrubia et al., 2008; Benson, 2012; Du et al., 2024).
@@
In this paper, we present a novel method to investigate the astrometric shifts of images in cluster lenses. The fundamental goal of this work is to constrain the CDM subhalo SHMF based on the induced perturbations from populations of CDM subhalos near the lensing critical curve. To ease comparison with previous works, we focus on the mean subhalo mass fraction ($f_{\rm sub}$) as the primary parameter of interest, implicitly assuming that it is the same in both galaxy and cluster environments. This parameter can be directly calculated from the subhalo SHMF, and has been constrained in many studies using galaxy-scale lenses (Vegetti et al., 2014; Despali and Vegetti, 2017; Hsueh et al., 2020; Gilman et al., 2020), but not with cluster-scale lenses. For our method, we utilize a known result from gravitational lensing in that sources that form near a fold caustic will form symmetric image pairs across the critical curve (Schneider et al., 1992). Thus, astrometric perturbations from a population of CDM subhalos lying in the lens plane will manifest as an asymmetry [^1] The induced asymmetry from CDM subhalos is akin to “Denting the mirror”, hence the title. in the image positions across the critical curve (Dai et al., 2018). We develop a methodology to infer the hidden subhalo population using a likelihood-free inference method that uses the asymmetry of lensed image pairs as a summary statistic. After validating the modeling framework on simulated datasets, we apply it to two lensed arcs in MACSJ0416 and AS1063 to infer the projected mass fraction in dark matter subhalos. In the future, the method demonstrated in this paper will be applied to larger samples of gravitational lenses to provide stringent constraints on $f_{\rm sub}$ from cluster lenses.
@@ 2.1 Gravitational Lensing Formalism
Here, we briefly review the strong gravitational lensing formalism necessary for this article. We refer the reader to various review articles (e.g Blandford and Narayan, 1986; Schneider et al., 1992; Narayan and Bartelmann, 1996) for additional details.
@@ 2.2 Lensing Near the Critical Curve
Cluster lens modeling analyses often model extended lensed arcs using identified counterimaged “knots” within the arc as constraints. Modeling the relative positions of lensed knots improves the cluster lens model by constraining the deflection field on angular scales comparable to the image separation (e.g. Bergamini et al., 2023; Perera et al., 2025b). These knots likely correspond to bright sub-galactic structures within the source galaxy (e.g. HII regions in the case of MACSJ1149 Williams et al., 2024a). Due to these knots forming on sub-arcsecond scales, they are treated as point images with lower positional uncertainty than a typical cluster scale image. For this paper, we represent lensed arcs as a collection of lensed point image knots across a critical curve.
@@
The prediction that image pairs collect along a straight line is a well known result in gravitational lensing near a fold caustic (Blandford and Narayan, 1986; Schneider et al., 1992). We derive this property below. We begin with the lensing Jacobian as defined in equation 5. It is more useful to write the Jacobian in terms of the convergence and shear tensor components ($\gamma_{1}$ and $\gamma_{2}$):
@@
Usefully, $\mathcal{A}(\boldsymbol{\theta})=\partial\boldsymbol{\beta}/\partial\boldsymbol{\theta}$. Since we are close to the critical curve, we can use the same approximations as before, namely that the origin is located at the critical curve position ($\boldsymbol{\theta_{c}}=0$). We define the components of the displacement $\delta\boldsymbol{\theta}=(\delta\theta_{1},\delta\theta_{2})$ and source position $\boldsymbol{\beta}=(\beta_{1},\beta_{2})$. The subscripts 1 and 2 indicate the tangential and critical directions in the principal coordinate system [^2] These subscripts are not the same as those of the shear tensor components. $\gamma_{1,0}$ in this coordinate system does not change to $\gamma_{2,0}$ when calculating $\beta_{2}$. Instead, the eigenvalue is what changes in this coordinate frame, where $1-\kappa-\gamma_{1,0}=0$. . The leading order terms can be written as:
@@
Since $\delta\theta_{2}$ is the critical direction (perpendicular to the critical curve), this shows that the formed images near the caustic fold are in fact symmetric on either side. As a consequence, the midpoints of these images will form at the critical curve. It should be noted that including higher order terms in equation 13 will cause the image midpoints to shift slightly off the critical curve. We demonstrate that this additional shift is negligible in most cases in Appendix B.
@@ 3.1.1 Cluster Lens Profile
To demonstrate the methodology presented in this work, we create a catalog of simulated cluster lens systems. These systems are comprised of a macro cluster lens profile, plus subhalos. We model the macro cluster-scale lens profile utilizing 3 main halos at a lens redshift $z_{d}=0.25$. Two of the halos are separated by 0.5" (1.97 kpc) such that their mass profile is effectively combined. This is done to mimic a slightly non-elliptical profile. The third profile is located a distance of 17.5" (68.8 kpc) from the other combined halos. Each halo is modelled as a Non-Singular Isothermal Ellipsoid (NSIE), which has an analytical lens potential (Hinshaw and Krauss, 1987; Halkola et al., 2006):
@@
where $\sigma_{v}$ is the velocity dispersion, $R_{c}$ is the angular core radius [^3] The physical core radius for the NSIE is equal to $D_{d}R_{c}$., $q$ is the axis ratio, and the vector position $\boldsymbol{\theta}$ has components $(\theta_{1},\theta_{2})$. The total mass of the cluster-scale lens profile is $1.3\times 10^{14}M_{\odot}$. Large clusters such as this are the most efficient lenses akin to the Frontier Fields (Lotz et al., 2017), motivating our lens profile as reasonable and realistic. Properties of each NSIE halo are presented in Table 1. Figure 2 shows the surface mass density of the cluster lens. From here, we refer to this profile as the fiducial macrolens.
@@ 3.1.2 Lensed Arcs Near Caustic Folds
In Section 2.2 we described the general properties of lensed images near the critical curve. We also showed that sources that form near the caustic folds of cluster-scale gravitational lenses will appear as fully symmetric lensed arcs across a linear critical curve for a smooth lens profile (Blandford and Narayan, 1986; Schneider et al., 1992; Dai et al., 2018). Here, we focus specifically on the corresponding approximations that we adopt.
@@
The addition of dark matter substructures in the region within $\sim 1-2$" of the critical curve will displace the smooth critical curve and can produce asymmetries in the arcs. It should be noted that the subhalo mass fraction is constrained to be on the order of $\sim 1\%$ (Gilman et al., 2019; Dai et al., 2020; Ji and Dai, 2025), but these constraints come mainly from galaxy-scale lens systems. The addition of dark matter subhalos will not cause lensing perturbations on scales beyond the $\sim 1-2$" window. Therefore, we only simulate subhalos within the immediate vicinity of lensed arcs.
@@
According to gravitational lensing theory, the midpoints of images that form near a smooth critical curve lie along the critical curve (Blandford and Narayan, 1986; Venumadhav et al., 2017; Dai et al., 2018) as we showed in Section 2.2. This symmetric setup can be broken with the inclusion of dark matter subhalos near the critical curve, as shown in Figure 1. Recall that images in these cases where the source is very near to the caustic manifest as extended arcs and are instead represented as lensed point image knots. Because of this, the image midpoints are easily calculated as the midpoints of the counterimaged knots. We note that this scenario where the midpoints lie along the critical curve is restricted to sources that form near to the caustic, and is not generally true for all source positions.
@@
When considering angular scales significantly smaller than the radius of curvature of the macrolens near the caustic fold, the forward lens-projected (i.e. mapped from the source to lens plane) critical curve can be approximated as a straight line, and we expect unperturbed image midpoints to collect along this straight line. In real examples of arcs, this appears to be relatively common (e.g. Kaurov et al., 2019; Chen et al., 2019; Diego et al., 2023). In fact, this is an intuitively expected setup for lensed arcs, given that in clusters the caustic fold spans a significantly larger region than cusps. It should also be noted that this assumption regarding the image midpoints is commonly used as a proxy for the true location of the smooth critical curve (Kelly et al., 2022; Broadhurst et al., 2025), emphasizing the utility of this approximation.
@@
It is important to note that there can be a degeneracy when evaluating lensed arcs along the critical curve. Deviations from a smooth fold can occur due to the presence of galaxy scale perturbers, such as cluster member galaxies (e.g. Dai et al., 2020). In this case, the midpoints will still lie along the critical curve, but the critical curve will no longer be a straight line. Likewise, large subhalos ($\gtrsim 10^{10}M_{\odot}$) may also contribute, albeit rarely, to the variance in the critical curve linearity. Furthermore, sources forming slightly away from the caustic fold will produce images across an intrinsically curved smooth critical curve, thus not allowing the straight line critical curve approximation. As such, there exists a degeneracy between deviations from the symmetry of arcs stemming from either dark matter subhalos or lens model complexities. For this work, we restrict the analysis to lensed arcs that are unaffected by cluster members or large subhalos and are subject to only the smooth macrolens near a caustic fold. It turns out that this need not be a strict requirement, but we caution anyway that our results should only be applied to lensed arcs that are consistent with this description. We examine in greater detail the limits of this assumption in Appendix A. In future works, these contributions away from linearity should be included in the statistical analysis.